an operator needs 400 gallons of a 5% solution mix, how many gallons of a 3% solution mix and an 8% solution mix does he need to mix
Answer by lwsshak3(11628)

an operator needs 400 gallons of a 5% solution mix, how many gallons of a 3% solution mix and an 8% solution mix does he need to mix
**
let x=amt of 3% solution to mix
400-x=amt of 8% solution to mix
3%x+8%(400-x)=5%*400
.03x+32-.08x=20
.05x=12
x=240
400-x=160
amt of 3% solution to mix=240 gal
amt of 8% solution to mix=160 gal